• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hoyoung

    Log entry time 07:59:12 on March 03, 2012

    Entry number 362940

    keyword=Owl Shift Summary March 03, 2012

    Shift workers: Hoyoung Kang(SL), Toby Badman(TO), Marco Pannunzio Carmignotto(3rd)

    Plan: BPM calibration

    00:20 BPM calibration, but HARP software has a problem 01:00 got 5uA for beam tuning 01:10 set BPMA to 0.0/0.0 to do HARP scan for nine points, HARP by hand 01:25 set BPMA to 5.0/0.0, but beam tripped for long time 02:20 Beam back as separator restores, re-do HARP scan for the position 02:45 100 nA CW beam with raster on, fast sets 2mm x 2mm and slow 1 cm 04:20 Controlled access requested to adjust BPM gains and to check fast raster signal 06:20 Work in hall finished, coming back to beam permit, but MCC still has problems 06:30 LHe was not filled up for pid loop error, so TO manually filled it 07:30 Beam is back. BPM calibration resummed. 07:55 requested 200nA pulse beam with raster off